# GRINNs: Godunov-Riemann Informed Neural Networks for Learning Hyperbolic Conservation Laws

### v1

@@ Abstract
We present GRINNs: numerical analysis-informed neural networks for the solution of inverse problems of non-linear systems of conservation laws. GRINNs are based on high-resolution Godunov schemes for the solution of the Riemann problem in hyperbolic Partial Differential Equations (PDEs). In contrast to other existing machine learning methods that learn the numerical fluxes of conservative Finite Volume methods, GRINNs learn the physical flux function per se. Due to their structure, GRINNs provide interpretable, conservative schemes, that learn the solution operator on the basis of approximate Riemann solvers that satisfy the Rankine-Hugoniot condition. The performance of GRINNs is assessed via four benchmark problems, namely the Burger’s, the Shallow Water, the Lighthill-Whitham-Richards and the Payne-Whitham traffic flow models. The solution profiles of these PDEs exhibit shock waves, rarefactions and/or contact discontinuities at finite times. We demonstrate that GRINNs provide a very high accuracy both in the smooth and discontinuous regions.
@@ 1 Introduction
In this work, we present a hybrid approach for the solution of the inverse problem for both scalar and systems of hyperbolic PDEs in 1-dim. spatial domains. In particular, we couple NNs with high-resolution Godunov-type FV schemes for the solution of the Riemann problem. We name this scheme as Godunov-Riemann Informed Neural Networks (GRINNs).
@@
In contrast to existing hybrid methods that discover numerical fluxes [54, 12, 36], here we use shallow NNs to approximate directly the physical flux functional terms, which are missing from the flux function of the unknown system of PDEs. To take into account partial information that may be available, GRINNs can be also employed in the presence of known fluxes or source terms in the system of PDEs. For obtaining the numerical solution of the unknown system, we perform operator splitting and solve the homogeneous part of the PDEs with high-resolution Godunov-type methods with flux-limiters and the inhomogeneous part (if any) with stiff integration schemes. These numerical schemes provide high-order accuracy both far and near the emerging discontinuities. However, since we deal with non-linear systems of hyperbolic PDEs, the related Riemann problem is solved by approximate Riemann solvers [41], which need to satisfy the Rankine-Hugoniot (RH) condition in order for the numerical scheme to be conservative [73, 41]. While NNs have been used to obtain accurate and RH condition-satisfying approximations for the solution of the Riemann problem via regression [49, 68, 76], here we solve the inverse problem with unknown flux terms; hence the target solutions are simply not available. To overcome this obstacle, we construct Riemann solvers that are dependent on the unknown flux function and constrain them to satisfy the RH condition; the latter is implemented as a soft constraint.
@@
In this way, GRINNs not only learn the physical unknown flux term/closures, but also the related approximate Riemann solver that renders the scheme conservative. This enables the direct implementation of GRINNs into software packages for the solution of hyperbolic PDEs, as Clawpack [14], which require both the functional and the Riemann solver. This is the first time that such a hybrid approach for systems of conservation laws, is presented, allowing also the inclusion of source terms.
@@
The performance of GRINNs is assessed via four benchmark problems of non-linear conservation laws arising in fluid dynamics and traffic flow. First, we consider the scalar conservation laws of the inviscid Burgers’ equation [9] and the Lighthill-Whitham-Richards (LWR) equation [44, 64], for which we learn the flux function and the velocity closure, respectively. Next, we study two systems of two conservation laws, the Shallow Water (SW) and the Payne-Whitham (PW) equations [59, 80], for both of which we learn the pressure closures in the flux function. We highlight here, that the Burgers’, LWR and SW equations are homogeneous hyperbolic PDEs that exhibit analytic solutions for the Riemann problem, while for the PW equations not only the Riemann problem cannot be solved analytically, but also source terms are included. For all the benchmark problems, the approximate Riemann solvers are derived in appendix C. We underline here that for all problems under study, GRINNs do not require extensive training data sets for achieving high approximation accuracy. This is because, a low number of complete (in space), time observations results in a high number of residuals to be minimized for the optimization problem, even for coarse spatial discretizations.
@@
The reminder of the paper is organized as follows. In section 2, we concisely describe the problem definition. In section 3, we briefly present the high-resolution Godunov-type FV numerical scheme and introduce in detail GRINNs for the solution of the inverse problem. In section 4, we present the numerical results for the four benchmark problems considered (which are presented in Appendix A along with details about the forward problem solution required for the data acquisition) and assess the convergence and accuracy of GRINNs. Finally, in section 5, we conclude, discussing limitations and perspectives of the present work.
@@ 2 Problem Definition
3 Methodology: The GRINNs
We present Godunov-Riemann Informed Neural Networks GRINNs, a hybrid numerical analysis informed SciML scheme, for solving the inverse problem for hyperbolic systems of PDEs in the form of eq. 2, particularly for learning unknown flux terms. Our approach relies on the solution of the forward problem via high-resolution Godunov-type FV schemes, which, for the sake of completeness, we briefly describe below.
@@ 3.1 High-resolution Godunov-type methods
3.2 Solution of the inverse problem with GRINNs
Here, we present GRINNs for the solution of the inverse problem for scalar or systems of conservation laws described by hyperbolic PDEs. As already discussed, this is a hybrid approach combining the use of NNs with the high-resolution Godunov-type numerical methods presented in Section 3.1.
Let us first revisit the system of hyperbolic PDEs in eq. 2 with the unknown flux functionals term $\mathbf{f}_{U}(\mathbf{u},\mathcal{N}(\mathbf{u}))$. For the approximation of the unknown functional, we consider a shallow NN, $\mathcal{N}(\mathbf{u}):=\mathcal{N}(\mathbf{u},\mathbf{P};\mathbf{H}_{1}):\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$, with network parameters $\mathbf{P}$ (weights and biases of each layer) and hyper-parameters $\mathbf{H}_{1}$ (e.g., activation function, learning rate, the number of epochs, etc.). To solve the inverse problem with GRINNs, we seek to find the parameters $\mathbf{P}$ of the NN $\mathcal{N}(\mathbf{u},\mathbf{P};\mathbf{H}_{1})$, such that the employment of the high-resolution Godunov-type method of eq. 6 to the system in eq. 2 generates the available data for the values of the state variables.
GRINNs as conservative schemes
@@ 3.2 Solution of the inverse problem with GRINNs
where the NN function is computed at $\bar{\mathbf{Q}}_{i}^{n}$, $\mathbf{Q}^{n}_{i}$ and $\mathbf{Q}^{n}_{i-1}$. Finding the networks parameters that satisfy eq. 13 ensures that the proposed GRINNs are conservative.
High-resolution GRINNs
@@
GRINNs optimization problem
@@
Algorithm 1 GRINNs implementation
@@
$\triangleright$ Procedure to compute the speeds and the waves provided by the approximate Riemann solver of the GRINNs, for given left and right states $\mathbf{q}_{l},\mathbf{q}_{r}$, NN parameters $\mathbf{P}$ and hyperparameters $\mathbf{H}_{1}$, $\mathbf{H}_{2}$
@@
GRINNs implementation
For the solution of the inverse problem with the GRINNs, we employed a single-layer feedforward NN architecture with $D$-dim. inputs and 1-dim. output, since for all benchmark problems examined, a scalar flux function term was assumed unknown. For $L$ neurons in the hidden layer, the output of the NN is written as:
@@
For the solution of the GRINNs optimization problem in eq. 15, the two loss functions in eqs. 16 and 17 were expressed in an $(D\times N\times n_{t})$-dim. and an $(N\times n_{t})$-dim. vector of non-linear residuals, say $\mathcal{F}_{FVG}(\mathbf{P})$ and $\mathcal{F}_{RH}(\mathbf{P})$, respectively. To minimize these residuals w.r.t. the unknown parameters $\mathbf{P}$, we used the Levenberg-Marquardt (LM) algorithm [28]. LM is a deterministic gradient-based optimization algorithm, which updates the unknown parameters according to a varying - at each iteration - damping factor $\lambda$. It thus, requires at each iteration the derivatives $\partial_{\mathbf{P}}\mathcal{F}_{FVG}$ and $\partial_{\mathbf{P}}\mathcal{F}_{RH}$. Since symbolic differentiation of $\partial_{\mathbf{P}}\mathcal{F}_{FVG}$ is not feasible, we used numerical differentiation through finite differences. Note that the residuals corresponding to eq. 16 can be computed independently for each $n=1,\ldots,n_{t}$, a feature that allowed us to perform parallel computations to form $\mathcal{F}_{FVG}(\mathbf{P})$ and thus accelerate convergence of the LM algorithm.
@@
We highlight here an important feature of GRINNs, related to the optimization problem and the LM algorithm employed for its solution. As discussed, GRINNs are specifically designed with a linearized Riemann solver that renders the scheme conservative and consistent with the unknown non-linear system of conservation laws in eq. 2. However, there is no guarantee that the linearized system is hyperbolic, i.e., that the GRINNs Roe matrix in eq. 11 is diagonalizable with real eigenvalues. To ensure hyperbolicity throughout the optimization process, one may include it as a hard constraint in the GRINNs optimization problem in eq. 15, the solution of which would then require constrained optimization algorithms. Here, to avoid such algorithms, we initialized the LM algorithm with a random guess of the parameters $\mathbf{P}$, such that the Roe matrix in eq. 11 is diagonalizable with real eigenvalues. For ensuring hyperbolicity throughout the optimization process with the LM algorithm, a high value of the initial damping factor should be avoided for averting big steps in the parameter space, which may lead to loss of hyperbolicity.
3.3 Implementation and numerical assessment of GRINNs
@@
In this section, we provide all the practical details for the implementation and numerical assessment of GRINNs, as employed for learning the unknown flux terms in the four benchmark problems considered.
@@
For acquiring the data upon which GRINNs are trained/tested on, we solved the forward problem for the known hyperbolic systems in the form of eq. 1 with well-posed initial data and boundary conditions. In particular, we consider two types of boundary conditions; either (i) periodic, in which the flow exits from the left/right boundary and enters from the right/left with the same characteristics, or (ii) outflow, in which the flow exits from the boundaries without reflections. For the former, we set the ghost cells to $\mathbf{Q}^{n}_{-1}=\mathbf{Q}^{n}_{N-1}$, $\mathbf{Q}^{n}_{0}=\mathbf{Q}^{n}_{N}$, $\mathbf{Q}^{n}_{N+1}=\mathbf{Q}^{n}_{1}$ and $\mathbf{Q}^{n}_{N+2}=\mathbf{Q}^{n}_{2}$ at every time step, while for the latter, we use zero-order extrapolation by setting $\mathbf{Q}^{n}_{-1}=\mathbf{Q}^{n}_{0}=\mathbf{Q}^{n}_{1}$ and $\mathbf{Q}^{n}_{N+1}=\mathbf{Q}^{n}_{N+2}=\mathbf{Q}^{n}_{N}$. For each problem, we consider smooth Gaussian or sinusoidal initial conditions $\mathbf{u}(0,x)$, which lead to shock waves and/or rarefactions along the time interval considered. For employing the high-resolution Godunov-type method described in Section 3.1, we compute the waves and speeds via the Roe linearized matrices derived for each problem, except from the fourth benchmark problem of the PW equations, for which the HLLE solver was used. In addition, in all problems, we used the Van-Leer flux-limiter function to derive the limited-versions of the waves in eq. 7. Finally, for the PW equations, where a source term is included in eq. 1, we used the stiff numerical integration scheme provided by ode15s of MATLAB ODE suite [70] for integrating the inhomogeneous part, as described in Section 3.1.
@@
For each benchmark problem, we solved the forward problem for $4$ different initial conditions in the $t\in[0,t_{end}]$ interval with different, for each problem, $dt$ satisfying the CFL condition. From the collected data, we randomly split the 15-15-70% of the data to form the training, validation and testing sets; for the PW equations the split was 7.5-7.5-85% because the CFL-satisfying $dt$ was smaller. The resulting sets consist of pairs of data points $\mathbf{Q}_{i}^{n},\mathbf{Q}_{i}^{n+1}\in\mathbb{R}^{D}$ for all cells $i=1,\ldots,N$ and for randomly selected time steps $n=1,\ldots,n_{t}$. Only the training set was used for solving the GRINNs optimization problem in eq. 15, which resulted in the formation of a total of $(D+1)\times N\times n_{t}$ residuals. We emphasize here that a larger training data set is not required, since only a few, complete time observations $n_{t}$ are enough for forming many residuals, even with coarse spatial discretizations. For each problem, we determined $n_{t}$ so that the ratio between the resulting number of residuals and the total number of trainable parameters in $\mathbf{P}$ is $\sim 5000$; the number of neurons was set to $L=5$, resulting to $5(D+2)$ trainable parameters.
@@
As already discussed, for the GRINNs optimization problem we used the LM algorithm over the training set, initialized with a random guess of parameters satisfying the hyperbolicity property of the Roe matrix in eq. 11 and an initial damping factor $\lambda_{0}=0.01$. The stopping criteria of the LM algorithm were (i) a maximum number of $500$ epochs, (ii) a relative loss function ($l^{2}$) tolerance of $10^{-9}$ and (iii) a validation error below $10^{-9}$ for 3 consecutive iterations; the latter criterion was evaluated over the validation set in a frequency of $20$ training epochs. To evaluate the convergence of the GRINNs optimization problem, we computed the maximum and mean $l^{1}$ errors $\lVert\hat{\mathbf{Q}}_{i}^{n+1}-\mathbf{Q}_{i}^{n+1}\rVert_{1}$ and the $MSE(\hat{\mathbf{Q}}_{i}^{n+1},\mathbf{Q}_{i}^{n+1})$, over all cells $i=1,\ldots,N$, time steps $n=1,\ldots,n_{t}$, and state variable dimensions $D$; $\hat{\mathbf{Q}}_{i}^{n+1}$ is the predicted value of the state variable at the next time step, as shown in eq. 14. An outline of the algorithm for the solution of the GRINNs optimization problem with the LM iterative scheme is provided in Algorithm 2.
Algorithm 2 Outline of the solution of the GRINNs optimization problem in eq. 15 with the LM algorithm
@@
Finally, to assess the numerical accuracy of GRINNs, we solve the forward problem using the learned flux function $\mathbf{f}_{U}(\mathbf{u},\mathcal{N}(\mathbf{u}))$ in eq. 2 and compare its solution $\hat{\mathbf{Q}}_{i}^{n+1}$ with the ground truth solution $\mathbf{Q}_{i}^{n+1}$ of the forward problem for the analytically known flux function in eq. 1. To quantify the numerical accuracy, we use the absolute errors $\lvert\hat{\mathbf{Q}}_{i}^{n+1}-\mathbf{Q}_{i}^{n+1}\rvert$ all over cells and time steps for all the components of the state variables $D$. Note that this error is not the same with the training/validation/testing errors, since the solution of the forward problem with the learned flux function $\mathbf{f}_{U}(\mathbf{u},\mathcal{N}(\mathbf{u}))$ also takes into account the cumulative error introduced by the numerical scheme.

### v2

# GoRINNs: Godunov-Riemann Informed Neural Networks for Learning Hyperbolic Conservation Laws

@@ Abstract
We present GoRINNs: numerical analysis-informed neural networks for the solution of inverse problems of non-linear systems of conservation laws. GoRINNs are based on high-resolution Godunov schemes for the solution of the Riemann problem in hyperbolic Partial Differential Equations (PDEs). In contrast to other existing machine learning methods that learn the numerical fluxes of conservative Finite Volume methods, GoRINNs learn the physical flux function per se. Due to their structure, GoRINNs provide interpretable, conservative schemes, that learn the solution operator on the basis of approximate Riemann solvers that satisfy the Rankine-Hugoniot condition. The performance of GoRINNs is assessed via four benchmark problems, namely the Burgers’, the Shallow Water, the Lighthill-Whitham-Richards and the Payne-Whitham traffic flow models. The solution profiles of these PDEs exhibit shock waves, rarefactions and/or contact discontinuities at finite times. We demonstrate that GoRINNs provide a very high accuracy both in the smooth and discontinuous regions.
@@ 1 Introduction
In this work, we present a hybrid approach for the solution of the inverse problem for both scalar and systems of hyperbolic PDEs in 1-dim. spatial domains. In particular, we couple NNs with high-resolution Godunov-type FV schemes for the solution of the Riemann problem. We name this scheme as Godunov-Riemann Informed Neural Networks (GoRINNs).
@@
In contrast to existing hybrid methods that discover numerical fluxes [54, 12, 36], here we use shallow NNs to approximate directly the physical flux functional terms, which are missing from the flux function of the unknown system of PDEs. To take into account partial information that may be available, GoRINNs can be also employed in the presence of known fluxes or source terms in the system of PDEs. For obtaining the numerical solution of the unknown system, we perform operator splitting and solve the homogeneous part of the PDEs with high-resolution Godunov-type methods with flux-limiters and the inhomogeneous part (if any) with stiff integration schemes. These numerical schemes provide high-order accuracy both far and near the emerging discontinuities. However, since we deal with non-linear systems of hyperbolic PDEs, the related Riemann problem is solved by approximate Riemann solvers [41], which need to satisfy the Rankine-Hugoniot (RH) condition in order for the numerical scheme to be conservative [73, 41]. While NNs have been used to obtain accurate and RH condition-satisfying approximations for the solution of the Riemann problem via regression [49, 68, 76], here we solve the inverse problem with unknown flux terms; hence the target solutions are simply not available. To overcome this obstacle, we construct Riemann solvers that are dependent on the unknown flux function and constrain them to satisfy the RH condition; the latter is implemented as a soft constraint.
@@
In this way, GoRINNs not only learn the physical unknown flux term/closures, but also the related approximate Riemann solver that renders the scheme conservative. This enables the direct implementation of GoRINNs into software packages for the solution of hyperbolic PDEs, as Clawpack [14], which require both the functional and the Riemann solver. This is the first time that such a hybrid approach for systems of conservation laws, is presented, allowing also the inclusion of source terms.
@@
The performance of GoRINNs is assessed via four benchmark problems of non-linear conservation laws arising in fluid dynamics and traffic flow. First, we consider the scalar conservation laws of the inviscid Burgers’ equation [9] and the Lighthill-Whitham-Richards (LWR) equation [44, 64], for which we learn the flux function and the velocity closure, respectively. Next, we study two systems of two conservation laws, the Shallow Water (SW) and the Payne-Whitham (PW) equations [59, 80], for both of which we learn the pressure closures in the flux function. We highlight here, that the Burgers’, LWR and SW equations are homogeneous hyperbolic PDEs that exhibit analytic solutions for the Riemann problem, while for the PW equations not only the Riemann problem cannot be solved analytically, but also source terms are included. For all the benchmark problems, the approximate Riemann solvers are derived in appendix C. We underline here that for all problems under study, GoRINNs do not require extensive training data sets for achieving high approximation accuracy. This is because, a low number of complete (in space), time observations results in a high number of residuals to be minimized for the optimization problem, even for coarse spatial discretizations.
@@
The reminder of the paper is organized as follows. In section 2, we concisely describe the problem definition. In section 3, we briefly present the high-resolution Godunov-type FV numerical scheme and introduce in detail GoRINNs for the solution of the inverse problem. In section 4, we present the numerical results for the four benchmark problems considered (which are presented in Appendix A along with details about the forward problem solution required for the data acquisition) and assess the convergence and accuracy of GoRINNs. Finally, in section 5, we conclude, discussing limitations and perspectives of the present work.
@@ 2 Problem Definition
3 Methodology: The GoRINNs
We present Godunov-Riemann Informed Neural Networks GoRINNs, a hybrid numerical analysis informed SciML scheme, for solving the inverse problem for hyperbolic systems of PDEs in the form of eq. 2, particularly for learning unknown flux terms. Our approach relies on the solution of the forward problem via high-resolution Godunov-type FV schemes, which, for the sake of completeness, we briefly describe below.
@@ 3.1 High-resolution Godunov-type methods
3.2 Solution of the inverse problem with GoRINNs
Here, we present GoRINNs for the solution of the inverse problem for scalar or systems of conservation laws described by hyperbolic PDEs. As already discussed, this is a hybrid approach combining the use of NNs with the high-resolution Godunov-type numerical methods presented in Section 3.1.
Let us first revisit the system of hyperbolic PDEs in eq. 2 with the unknown flux functionals term $\mathbf{f}_{U}(\mathbf{u},\mathcal{N}(\mathbf{u}))$. For the approximation of the unknown functional, we consider a shallow NN, $\mathcal{N}(\mathbf{u}):=\mathcal{N}(\mathbf{u},\mathbf{P};\mathbf{H}_{1}):\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$, with network parameters $\mathbf{P}$ (weights and biases of each layer) and hyper-parameters $\mathbf{H}_{1}$ (e.g., activation function, learning rate, the number of epochs, etc.). To solve the inverse problem with GoRINNs, we seek to find the parameters $\mathbf{P}$ of the NN $\mathcal{N}(\mathbf{u},\mathbf{P};\mathbf{H}_{1})$, such that the employment of the high-resolution Godunov-type method of eq. 6 to the system in eq. 2 generates the available data for the values of the state variables.
GoRINNs as conservative schemes
@@ 3.2 Solution of the inverse problem with GoRINNs
where the NN function is computed at $\bar{\mathbf{Q}}_{i}^{n}$, $\mathbf{Q}^{n}_{i}$ and $\mathbf{Q}^{n}_{i-1}$. Finding the networks parameters that satisfy eq. 13 ensures that the proposed GoRINNs are conservative.
High-resolution GoRINNs
@@
GoRINNs optimization problem
@@
Algorithm 1 GoRINNs implementation
@@
$\triangleright$ Procedure to compute the speeds and the waves provided by the approximate Riemann solver of the GoRINNs, for given left and right states $\mathbf{q}_{l},\mathbf{q}_{r}$, NN parameters $\mathbf{P}$ and hyperparameters $\mathbf{H}_{1}$, $\mathbf{H}_{2}$
@@
GoRINNs implementation
For the solution of the inverse problem with the GoRINNs, we employed a single-layer feedforward NN architecture with $D$-dim. inputs and 1-dim. output, since for all benchmark problems examined, a scalar flux function term was assumed unknown. For $L$ neurons in the hidden layer, the output of the NN is written as:
@@
For the solution of the GoRINNs optimization problem in eq. 15, the two loss functions in eqs. 16 and 17 were expressed in an $(D\times N\times n_{t})$-dim. and an $(N\times n_{t})$-dim. vector of non-linear residuals, say $\mathcal{F}_{FVG}(\mathbf{P})$ and $\mathcal{F}_{RH}(\mathbf{P})$, respectively. To minimize these residuals w.r.t. the unknown parameters $\mathbf{P}$, we used the Levenberg-Marquardt (LM) algorithm [28]. LM is a deterministic gradient-based optimization algorithm, which updates the unknown parameters according to a varying - at each iteration - damping factor $\lambda$. It thus, requires at each iteration the derivatives $\partial_{\mathbf{P}}\mathcal{F}_{FVG}$ and $\partial_{\mathbf{P}}\mathcal{F}_{RH}$. Since symbolic differentiation of $\partial_{\mathbf{P}}\mathcal{F}_{FVG}$ is not feasible, we used numerical differentiation through finite differences. Note that the residuals corresponding to eq. 16 can be computed independently for each $n=1,\ldots,n_{t}$, a feature that allowed us to perform parallel computations to form $\mathcal{F}_{FVG}(\mathbf{P})$ and thus accelerate convergence of the LM algorithm.
@@
We highlight here an important feature of GoRINNs, related to the optimization problem and the LM algorithm employed for its solution. As discussed, GoRINNs are specifically designed with a linearized Riemann solver that renders the scheme conservative and consistent with the unknown non-linear system of conservation laws in eq. 2. However, there is no guarantee that the linearized system is hyperbolic, i.e., that the GoRINNs Roe matrix in eq. 11 is diagonalizable with real eigenvalues. To ensure hyperbolicity throughout the optimization process, one may include it as a hard constraint in the GoRINNs optimization problem in eq. 15, the solution of which would then require constrained optimization algorithms. Here, to avoid such algorithms, we initialized the LM algorithm with a random guess of the parameters $\mathbf{P}$, such that the Roe matrix in eq. 11 is diagonalizable with real eigenvalues. For ensuring hyperbolicity throughout the optimization process with the LM algorithm, a high value of the initial damping factor should be avoided for averting big steps in the parameter space, which may lead to loss of hyperbolicity.
3.3 Implementation and numerical assessment of GoRINNs
@@
In this section, we provide all the practical details for the implementation and numerical assessment of GoRINNs, as employed for learning the unknown flux terms in the four benchmark problems considered.
@@
For acquiring the data upon which GoRINNs are trained/tested on, we solved the forward problem for the known hyperbolic systems in the form of eq. 1 with well-posed initial data and boundary conditions. In particular, we consider two types of boundary conditions; either (i) periodic, in which the flow exits from the left/right boundary and enters from the right/left with the same characteristics, or (ii) outflow, in which the flow exits from the boundaries without reflections. For the former, we set the ghost cells to $\mathbf{Q}^{n}_{-1}=\mathbf{Q}^{n}_{N-1}$, $\mathbf{Q}^{n}_{0}=\mathbf{Q}^{n}_{N}$, $\mathbf{Q}^{n}_{N+1}=\mathbf{Q}^{n}_{1}$ and $\mathbf{Q}^{n}_{N+2}=\mathbf{Q}^{n}_{2}$ at every time step, while for the latter, we use zero-order extrapolation by setting $\mathbf{Q}^{n}_{-1}=\mathbf{Q}^{n}_{0}=\mathbf{Q}^{n}_{1}$ and $\mathbf{Q}^{n}_{N+1}=\mathbf{Q}^{n}_{N+2}=\mathbf{Q}^{n}_{N}$. For each problem, we consider smooth Gaussian or sinusoidal initial conditions $\mathbf{u}(0,x)$, which lead to shock waves and/or rarefactions along the time interval considered. For employing the high-resolution Godunov-type method described in Section 3.1, we compute the waves and speeds via the Roe linearized matrices derived for each problem, except from the fourth benchmark problem of the PW equations, for which the HLLE solver was used. In addition, in all problems, we used the Van-Leer flux-limiter function to derive the limited-versions of the waves in eq. 7. Finally, for the PW equations, where a source term is included in eq. 1, we used the stiff numerical integration scheme provided by ode15s of MATLAB ODE suite [70] for integrating the inhomogeneous part, as described in Section 3.1.
@@
For each benchmark problem, we solved the forward problem for $4$ different initial conditions in the $t\in[0,t_{end}]$ interval with different, for each problem, $dt$ satisfying the CFL condition. From the collected data, we randomly split the 15-15-70% of the data to form the training, validation and testing sets; for the PW equations the split was 7.5-7.5-85% because the CFL-satisfying $dt$ was smaller. The resulting sets consist of pairs of data points $\mathbf{Q}_{i}^{n},\mathbf{Q}_{i}^{n+1}\in\mathbb{R}^{D}$ for all cells $i=1,\ldots,N$ and for randomly selected time steps $n=1,\ldots,n_{t}$. Only the training set was used for solving the GoRINNs optimization problem in eq. 15, which resulted in the formation of a total of $(D+1)\times N\times n_{t}$ residuals. We emphasize here that a larger training data set is not required, since only a few, complete time observations $n_{t}$ are enough for forming many residuals, even with coarse spatial discretizations. For each problem, we determined $n_{t}$ so that the ratio between the resulting number of residuals and the total number of trainable parameters in $\mathbf{P}$ is $\sim 5000$; the number of neurons was set to $L=5$, resulting to $5(D+2)$ trainable parameters.
@@
As already discussed, for the GoRINNs optimization problem we used the LM algorithm over the training set, initialized with a random guess of parameters satisfying the hyperbolicity property of the Roe matrix in eq. 11 and an initial damping factor $\lambda_{0}=0.01$. The stopping criteria of the LM algorithm were (i) a maximum number of $500$ epochs, (ii) a relative loss function ($l^{2}$) tolerance of $10^{-9}$ and (iii) a validation error below $10^{-9}$ for 3 consecutive iterations; the latter criterion was evaluated over the validation set in a frequency of $20$ training epochs. To evaluate the convergence of the GoRINNs optimization problem, we computed the maximum and mean $l^{1}$ errors $\lVert\hat{\mathbf{Q}}_{i}^{n+1}-\mathbf{Q}_{i}^{n+1}\rVert_{1}$ and the $MSE(\hat{\mathbf{Q}}_{i}^{n+1},\mathbf{Q}_{i}^{n+1})$, over all cells $i=1,\ldots,N$, time steps $n=1,\ldots,n_{t}$, and state variable dimensions $D$; $\hat{\mathbf{Q}}_{i}^{n+1}$ is the predicted value of the state variable at the next time step, as shown in eq. 14. An outline of the algorithm for the solution of the GoRINNs optimization problem with the LM iterative scheme is provided in Algorithm 2.
Algorithm 2 Outline of the solution of the GoRINNs optimization problem in eq. 15 with the LM algorithm
@@
Finally, to assess the numerical accuracy of GoRINNs, we solve the forward problem using the learned flux function $\mathbf{f}_{U}(\mathbf{u},\mathcal{N}(\mathbf{u}))$ in eq. 2 and compare its solution $\hat{\mathbf{Q}}_{i}^{n+1}$ with the ground truth solution $\mathbf{Q}_{i}^{n+1}$ of the forward problem for the analytically known flux function in eq. 1. To quantify the numerical accuracy, we use the absolute errors $\lvert\hat{\mathbf{Q}}_{i}^{n+1}-\mathbf{Q}_{i}^{n+1}\rvert$ all over cells and time steps for all the components of the state variables $D$. Note that this error is not the same with the training/validation/testing errors, since the solution of the forward problem with the learned flux function $\mathbf{f}_{U}(\mathbf{u},\mathcal{N}(\mathbf{u}))$ also takes into account the cumulative error introduced by the numerical scheme.